Northwood Company manufactures basketballs. The company has a ball that sells for $25. At present, the ball is manufactured in a small plant that relies heavily on direct labor workers. Thus, variable expenses are high, totaling $15.00 per ball, of which 60% is direct labor cost.

Last year, the company sold 38,000 of these balls, with the following results:

Sales (38,000 balls) $ 1,150,000
Variable expenses 690,000
Contribution margin 460,000
Fixed expenses 242,000
Net operating income $ 218,000

5. Refer to the original data. The company is discussing the construction of a new, automated manufacturing plant. The new plant would slash variable expenses per ball by 40.00%, but it would cause fixed expenses per year to double. If the new plant is built, what would be the company’s new CM ratio and new break-even point in balls?

6. Refer to the data in (5) above.

b. Assume the new plant is built and that next year the company manufactures and sells 38,000 balls (the same number as sold last year). Prepare a contribution format income statement and compute the degree of operating leverage.
Answer #1

Answer 5.

New variable costs = $690,000 * 60 % = $414,000

New fixed costs = $242,000 * 2 = $484,000

New contribution margin = Sales - New variable costs = $1,150,000 -  $414,000 = $736,000

Contribution margin per unit =$736,000 / 38,000 balls = $19.3684 per unit

Company’s new CM ratio = New contribution margin / Sales = $736,000 / $1,150,000 = 64 %

New break-even point in balls = New fixed costs / Contribution margin per unit

= $484,000 /$19.3684 per unit = 24,989 balls

Answer 6.

Contribution format income statement

Particular Amount ($)
Sales 1,150,000
- Variable expense (414,000)
Contribution margin 736,000
- Fixed expense (484,000)
Net income $252,000

Degree of operating leverage =  Contribution margin / Net income = $736,000 / $252,00 = 2.92